OpenStudy (anonymous):

I want solution of this question Check the continuity of f(x) at x=3 for the function defined as f(x)={3x+5 if x<3 f(x)=14 if x=3 f(x)=7+21/x if x>3

OpenStudy (agreene):

well, we want to determine if this is going to give us a hole/jump. So, lets check what happens just above and just below the point: x= 2, 3, 4 For f(2) = 3(2)+5 = 9 f(3) = 14 f(4) = 7+ 21/4 = 12.25 Well, it doesnt seem like much of a difference, let's see what all of these eqns do at x=3 3(3)+5 = 14 f(3) = 14 7+21/3 = 14 Looks like this is definitely continuous.

OpenStudy (anonymous):

The function is continuous. By definition, a function is continuous if the limit exists and are the same at a point approaching from both directions (i.e. from below or the left and above or the right), and the evaluated function is the same as the limit at that point.

OpenStudy (anonymous):

for aamir9119....agreena have answer our question...so juz follow ok